OpenStudy (anonymous):

Suppose you divide a polynomial by a binomial. How do you know if the binomial is a factor of the polynomial? Create a sample problem that has a binomial which IS a factor of the polynomial being divided, and another problem that has a binomial which is NOT a factor of the polynomial being divided.

OpenStudy (anonymous):

I dont understand where to start :/

OpenStudy (anonymous):

@TwilightRain1

OpenStudy (anonymous):

uhhhh i learned this a few months ago, but i forgot...

OpenStudy (anonymous):

its all good

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

@KendrickLamar2014

OpenStudy (anonymous):

@baller398

OpenStudy (anonymous):

alright, I'm here

OpenStudy (anonymous):

do you know how to do this sort of stuff

OpenStudy (anonymous):

to day is your lucky day yes i do so, After you've done the division, you know the binomial is a factor if the remainder is zero. (3x^3 - 2x^2 - 2x - 5) / (3x - 5) = x^2 + x + 1 Since there is no remainder, 3x - 5 is a factor of 3x^3 - 2x^2 - 2x - 5. (3x^3 - 2x^2 - 2x - 7) / (3x - 5) = x^2 + x + 1 with remainder -2. Since the remainder is not zero, 3x - 5 is not a factor of 3x^3 - 2x^2 - 2x - 7.

OpenStudy (anonymous):

tell me that helped!!!!

OpenStudy (anonymous):

it did, but I'm just horrible at algebra. i failed last year once we got into this sort of stuff dealing with binomials and all of that. do you mind explaining the division part? where you said "after you've done the division"

OpenStudy (anonymous):

@lsherron

OpenStudy (anonymous):

long division

OpenStudy (anonymous):

ill be right back i need to help som one realy qiick

OpenStudy (anonymous):

alright its all good

OpenStudy (anonymous):

I'm just trying to wrap my head around the concept. i think i understand it for the most part. I'm just bad with numbers. ill try to make my own problem now. one sec

OpenStudy (anonymous):

what numbers do you divide? to then get the remainder

OpenStudy (anonymous):

this will help

OpenStudy (anonymous):

ok ill check it out, thanks

OpenStudy (anonymous):

np

OpenStudy (anonymous):

i don't fully understand still

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@paki

OpenStudy (jackofalltradez):

@baller398

OpenStudy (sissyedgar):

lol, HERE!

OpenStudy (anonymous):

:]